OpenStudy (anonymous):

write a quadratic equation having the given numbers as solutions -7 and -2

OpenStudy (saifoo.khan):

x^2 + 9x + 14

OpenStudy (mertsj):

[(x-(-7)][x-(-2)]=0 (x+7)(x+2)=0 x^2+9x+14=0